# Scalable Data Synthesis for Computer Use Agents with Step-Level Filtering

### v1

@@ 5 Related Works
Computer use agents (CUAs) have recently emerged as a powerful class of autonomous systems that can operate directly on graphical user interfaces (operator; Claude-computer-use; agasheagent; qin2025ui; wang2025ui; wang2025opencua). Training these agents is particularly challenging because it requires grounding visual input into actionable affordances, reasoning and planning over long horizons, and executing precise actions in dynamic environments. While direct online interaction with GUIs would be the most natural way to acquire these capabilities, it is prohibitively expensive due to the need for virtual machines and complex infrastructure. As a result, most prior work relies on offline training with pre-collected trajectories.
Human-annotated datasets provide valuable supervision but are expensive and limited in scale. UI-TARS (qin2025ui) collected human demonstrations along with curated self-reflection examples, and UI-TARS-2 (wang2025ui) went further by requiring annotators to verbalize their reasoning traces during trajectory collection. Others, including MM-Mind2Web (zheng2024gpt) and GUIAct (chen2024guicourse), provide only short action trajectories without reasoning traces or step-level grading, while AgentTrek (xuagenttrek) includes reasoning but lacks explicit step quality control. Earlier datasets (liu2018reinforcement; yao2022webshop; koh2024visualwebarena) rely on Set-of-Marks or accessibility trees, which simplifies grounding but deviates from realistic deployment settings where only screenshots are available. These limitations motivate the need for scalable, reasoning-rich, and coordinate-based CUA data.
@@
With large-scale but noisy synthetic rollouts, filtering becomes crucial for ensuring that models learn from high-quality supervision. The dominant strategy in CUAs has been trajectory-level filtering, which retains only trajectories that reach successful task completion (sun-etal-2025-os). While straightforward, this approach is mismatched to the step-wise nature of CUA training, as successful trajectories often include suboptimal or even incorrect intermediate steps.
Broader works in reinforcement learning have explored step-level filtering, most notably filtered behavior cloning (FBC) (gulcehre2023reinforced; singhbeyond; shao2024deepseekmath), where the model selectively trains on high-quality actions. However, in CUAs, step-level filtering is especially challenging because rewards are inherently sparse: only final success is observed, and Monte Carlo credit assignment is high-variance. Actor-critic methods attempt to address this by learning a critic to provide intermediate feedback, as seen in DigiRL (bai2024digirl), but these methods are often unstable and costly to train (parisi2019td).
@@ 6 Conclusions
In this work, we present a scalable data synthesis pipeline for training computer use agents, addressing the limitations of costly human annotation and noisy trajectory-level filtering. Our key insight is that step-level filtering better aligns with the supervision objective, enabling agents to learn from correct actions even within imperfect trajectories. Combined with thought augmentation, which strengthens planning and stability, this approach yields substantial improvements across multiple benchmarks, surpassing strong baselines with only SFT. To support broader research, we introduce two complementary datasets: WebSTAR, a trajectory dataset enriched with reasoning and step-level scores, and WebSCORE, a grading dataset tailored for training reward models. Building on WebSCORE, we introduced StepRM, a lightweight reward model that matches o4-mini in grading quality at a lower cost, offering a practical tool for future reinforcement learning based optimization. Together, these contributions establish step-level filtering, reasoning-rich augmentation, and efficient reward modeling as core principles for robust and scalable CUA training.
@@
Domain All steps Correct steps UI-TARS-1.5-7B OAI CUA
Allrecipes 48.9 (80.0) 66.7 (86.7) 30.9 (56.8) 78.2 (91.1)
Amazon 29.3 (46.3) 31.7 (56.1) 33.5 (65.9) 54.3 (92.7)
Apple 25.6 (46.5) 41.9 (55.8) 33.7 (60.5) 68.6 (79.1)
Arxiv 20.9 (34.9) 41.9 (58.1) 20.9 (39.5) 71.5 (90.7)
BBCNews 38.1 (61.9) 64.3 (73.8) 38.1 (76.2) 67.1 (81.0)
Coursera 45.2 (66.7) 69.0 (78.6) 54.8 (76.2) 79.8 (92.9)
ESPN 27.3 (47.7) 25.0 (50.0) 23.3 (43.2) 56.8 (75.0)
GitHub 21.9 (75.6) 43.9 (82.9) 25.6 (48.8) 66.5 (85.4)
Huggingface 18.6 (25.6) 25.6 (34.9) 5.8 (11.6) 61.6 (79.1)
Wolfram 41.3 (65.2) 60.0 (65.2) 33.7 (47.8) 85.3 (93.5)
Average 31.9 (55.1) 47.0 (64.2) 30.0 (52.4) 69.2 (86.0)
Table 9: Performance comparison on 7B models. The reported metrics are Pass@1 and Pass@4 (in parentheses). The bold numbers show the best performing metric among the 7B models, where OAI CUA serves as the oracle. Step-level filtering (“Correct steps”) substantially improves over trajectory-level filtering (“All steps”), and further outperforms UI-TARS-1.5-7B across domains. OpenAI CUA (OAI CUA) remains strongest, but our method narrows the gap significantly.

### v2

@@ 5 Related Works
Computer use agents (CUAs) have recently emerged as a powerful class of autonomous systems that can operate directly on graphical user interfaces (OpenAI, 2025; Anthropic, 2024; Agashe et al., 2025; Qin et al., 2025; Wang et al., 2025a, b). Training these agents is particularly challenging because it requires grounding visual input into actionable affordances, reasoning and planning over long horizons, and executing precise actions in dynamic environments. While direct online interaction with GUIs would be the most natural way to acquire these capabilities, it is prohibitively expensive due to the need for virtual machines and complex infrastructure. As a result, most prior work relies on offline training with pre-collected trajectories.
Human-annotated datasets provide valuable supervision but are expensive and limited in scale. UI-TARS (Qin et al., 2025) collected human demonstrations along with curated self-reflection examples, and UI-TARS-2 (Wang et al., 2025a) went further by requiring annotators to verbalize their reasoning traces during trajectory collection. Others, including MM-Mind2Web (Zheng et al., 2024a) and GUIAct (Chen et al., 2024), provide only short action trajectories without reasoning traces or step-level grading, while AgentTrek (Xu et al., 2025a) includes reasoning but lacks explicit step quality control. Earlier datasets (Liu et al., 2018; Yao et al., 2022; Koh et al., 2024) rely on Set-of-Marks or accessibility trees, which simplifies grounding but deviates from realistic deployment settings where only screenshots are available. These limitations motivate the need for scalable, reasoning-rich, and coordinate-based CUA data.
@@
With large-scale but noisy synthetic rollouts, filtering becomes crucial for ensuring that models learn from high-quality supervision. The dominant strategy in CUAs has been trajectory-level filtering, which retains only trajectories that reach successful task completion (Sun et al., 2025). While straightforward, this approach is mismatched to the step-wise nature of CUA training, as successful trajectories often include suboptimal or even incorrect intermediate steps.
Broader works in reinforcement learning have explored step-level filtering, most notably filtered behavior cloning (FBC) (Gulcehre et al., 2023; Singh et al., ; Shao et al., 2024), where the model selectively trains on high-quality actions. However, in CUAs, step-level filtering is especially challenging because rewards are inherently sparse: only final success is observed, and Monte Carlo credit assignment is high-variance. Actor-critic methods attempt to address this by learning a critic to provide intermediate feedback (Bai et al., 2024, 2025a), but these methods are often unstable and costly to train (Parisi et al., 2019).
@@ 6 Conclusions
In this work, we present a scalable data synthesis pipeline for training computer use agents, addressing the limitations of costly human annotation and noisy trajectory-level filtering. Our key insight is that step-level filtering better aligns with the supervision objective, enabling agents to learn from correct actions even within imperfect trajectories. Combined with thought augmentation, which strengthens planning and stability, this approach yields substantial improvements across multiple benchmarks, surpassing strong baselines with only SFT. To support broader research, we introduce two complementary datasets: WebSTAR, a trajectory dataset enriched with reasoning and step-level scores, and WebSCORE, a grading dataset tailored for training process reward models. Building on WebSCORE, we introduced StepRM, a lightweight process reward model that matches o4-mini in grading quality at a lower cost, offering a practical tool for future reinforcement learning based optimization. Together, these contributions establish step-level filtering, reasoning-rich augmentation, and efficient reward modeling as core principles for robust and scalable CUA training.